A hollow cylindrical shaft has an internal diameter of 54 mm and a wall thickness of 4.7 mm. Calculate its polar second moment of area, J. Input your answer in units of 10^-9 m^4 to 2 decimal places. So, for example, if J is found to be 530.6x10^-12 m^4, you should input 0.53 as your answer.
Answer #1

Calculate the outer diameter as follows: d. = d; +2(t) = 54 +(2x4.7) d. = 63.4 mm = 63.4x10-3 m d: = 54x10-3 m Calculate the